OpenStudy (anonymous):

6x-3y =11 6x+3y =17 show work please

OpenStudy (curry):

6x-3y=11 + 6x + 3y= 17

OpenStudy (curry):

12x= 28

OpenStudy (curry):

28/12

OpenStudy (curry):

2 1/3 = x

OpenStudy (curry):

dont froget medal

OpenStudy (anonymous):

add both equations you will get, 12x=28 x=28/12 x=7/3 substitute x value in one of the equation 6(7/3)+3y =17 14+3y=17 3y=17-14 y=3/3 y=1 so,x=7/3 and y=1